15.0kg ball is dropped from 5.00m above ground. What is the speed of the ball just as it hits the ground?


1
Expert's answer
2021-04-28T09:56:13-0400

According to the conservation energy law, the kinetic energy of the ball just as it hits the ground is equal to its initial potential energy:


"E_p = E_k"

The potential energy can be calculated as follows:


"E_p = mgh"

where "m = 13kg" is the mass of the ball, "g = 9.8m\/s^2" is the gravitational acceleration, and "h = 5m" is the initial height.

The kinetic energy can be calculated as follows:


"E_k = \dfrac{mv^2}{2}"

where "v" is the required speed of the ball just as it hits the ground. Thus, obtain:


"mgh = \dfrac{mv^2}{2}\\
v = \sqrt{2gh}\\
v = \sqrt{2\cdot 9.8\cdot 5} \approx 9.9 m/s"

Answer. 9.9 m/s.